Let X1, X2 ... be a sequence of random variables. Write down the definition of Xin con- verges to X in distribution. State
Answer #1

4) (a) X1 , X2 , ..... sequence of RVs

Xn converges to X in distribution :

Notation : Xn  \overset{D}{\rightarrow} X

Definition : Consider a sequence {Xn} of RVs defined on probability space (\Omega , A, P) , with corresponding sequence of DFs {F_{n}(x)} . The sequence {Xn} of RVs converges in distribution/law to X , with distribution fn F , if  F_{n}(x) \rightarrow F(x) as n 1 00 , at every continuity points of F(x).

(b) Central Limit Theorem (by Lindeberg-levy): Let X1, ...Xn,... be a sequence of independent and identical RVs. \bar{X_{n}} = \frac{1}{n}\sum_{1=1}^{n} X_{i}

Then, \frac{\sqrt{n}(\bar{X_{n}-\mu})}{\sigma } \sim ^{a} N(0,1) as n \rightarrow \infty

where, \mu = E(X_{i}) ,\forall i & \sigma ^{2} = V(X_{i}) \forall i

(c) Let X = amount of time (in minutes) that each customer takes to pay

here, X\sim U([1,3])

\therefore \mu = E(X) = \frac{1+3}{2}= 2

\sigma ^{2} = V(X)=\frac{(3-1)^{2}}{12} = \frac{1}{3}

So, by CLT, \frac{\sqrt{n(\bar{X}-\mu)}}{\sigma } \sim ^{a} N(0,1) where Xi\sim ^{i\phi } U ([1,3])

Now probability that it will take less than 22 minutes in total for 10 customers to pay for their lines

= P\left [ \sum_{i=1}^{10}X_{i} < 22 \right ] = P\left [ \bar{X}< 2.2 \right ]

= P(\frac{\sqrt{10}(\bar{X}-2)}{\frac{1}{3}}<\frac{\sqrt{10}(2.2-2)}{\frac{1}{3}})

= P\left [ Z < 1.897 \right ] , where Z \sim N(0,1)

= \phi (1.897)= 0.9706